OpenStudy (anonymous):

Decide whether the equation is always, sometimes, or never true. 3 - 4x = 4x - 3

OpenStudy (anonymous):

Never true

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

how do you figure that out?

OpenStudy (anonymous):

Take some example (values) of x and try.

OpenStudy (anonymous):

but I set both to zero 3-4x=0 or 4x-3=0 -4x=-3 or 4x=3 x=3/4 or x=3/4 what did i do wrong?

OpenStudy (anonymous):

Nope I didn't mean that, we have : 3 - 4x = 4x - 3 // replace for example x = 0 3 - 4 (0) = 4 (0) - 3 3 = -3 // false And so on..